OpenStudy (anonymous):

16 OVER B= 2 OVER3? SHOW WORK

OpenStudy (saifoo.khan):

Cross muliply.

OpenStudy (saifoo.khan):

multiply*

OpenStudy (saifoo.khan):

b*2 = 16 * 3 b = 48/2 b = 24

OpenStudy (callisto):

16/B = 2/3 B= 16x3/2 = 24

OpenStudy (anonymous):

16/B= 2/3 or, B=16*(3/2) or, B=16*2.5 or, B=40